If I have a 62.38% in a class and my final is worth 40% of my grade, what grade do i need to get to pass? and if i got a 0, what will my final grade be?
Answer by jsmallt9(3758)

If the final is worth 40% of the final grade, then the 62.38% must be 60% of your final grade. So:
Final average = 0.60 * 62.38 + 0.4 * (final exam grade)

Since I do not know what grade is required to pass, I cannot specifically help with this part. I can only say that the minimum passing grade (60, 70, 75, ?) would go on the left side. Let's say it's 70. Then the equation becomes:
70 = 0.60 * 62.38 + 0.4 * (final exam grade)
Then use algebra to solve for the final exam grade to find out the lowest it could be in order to pass.

For the other part of the equation, just put a zero in for the final exam grade:
Final average = 0.60 * 62.38 + 0.4 * (0)
Simplifying...
Final average = 0.60 * 62.38 + 0
Final average = 0.60 * 62.38
Final average = 37.428
So if you get a zero on the final, your final average will be 37.428.